A disk between vertebrae in the spine is subjected to a shearing force of 716 N. Find its shear deformation, using the shear modulus of 1.0 × 109 N/m2. The disk is equivalent to a solid cylinder 0.63 cm high and 6.7 cm in diameter


1
Expert's answer
2021-03-29T09:09:56-0400
"S=\frac{FL}{Ax}\\x=\frac{FL}{AS}\\x=\frac{4FL}{\pi d^2S}
\\x=\frac{4(716)(0.63\cdot10^{-2})}{\pi(1\cdot10^{9})(6.7\cdot10^{-2})^2}=1.28\cdot10^{-6}\ m"